# The Origin Symphony: Probing Baryogenesis with Gravitational Waves

### v1

@@ Abstract
Affleck-Dine (AD) baryogenesis is compelling yet challenging to probe because of the high energy physics involved. We demonstrate that this mechanism can be realized generically with low-energy new physics without supersymmetry while producing detectable gravitational waves (GWs) sourced by parametric resonance of a light scalar field. In viable benchmark models, the scalar has a mass of ${\cal O}(0.1-10)$ GeV, yielding GWs with peak frequencies of ${\cal O}(10-100)$ Hz. This study further reveals a new complementarity between upcoming LIGO-frequency GW detectors and laboratory searches across multiple frontiers of particle physics.
Introduction.
@@
The origin of the observed matter-antimatter asymmetry $\eta=\frac{n_{b}-n_{\bar{b}}}{n_{\gamma}}\approx 6\times 10^{-10}$ Aghanim et al. (2020) (in terms of baryon-to-photon ratio) remains a long-standing puzzle in cosmology and particle physics, requiring new physics beyond the Standard Model (SM) to resolve it. Among the solutions proposed for the baryon asymmetry puzzle, Affleck-Dine (AD) baryogenesis Affleck and Dine (1985) is a compelling mechanism where the baryon asymmetry originates from the oscillation and subsequent decay of a baryon-number carrying scalar field. In its original form, the AD mechanism is constructed within the framework of supersymmetry (SUSY) Affleck and Dine (1985); Dine et al. (1996); Dine and Kusenko (2003); Enqvist and Mazumdar (2003), where the scalar condensates, potentially essential for baryogenesis, are flat direction moduli fields that are composed of superpartners of the SM quarks/leptons and the Higgs field. These flat direction moduli can generally acquire a large field value displacement during inflation, which spontaneously violates C and CP symmetries, then roll down, oscillate, and decay to baryons in the post-inflationary era. In conventional AD, the scalar field starts to roll and oscillate shortly after the end of inflation, while the inflaton condensate still dominates the Universe before its depletion via reheating Allahverdi and Mazumdar (2012). In alternative models, the AD field itself plays the role of the inflaton Hertzberg and Karouby (2014); Cline et al. (2020a).
Despite the attraction of the AD mechanism, its experimental test is generally challenging due to the high energy scale of the physics involved and the tie to the inflationary era. Fortunately, stochastic gravitational wave background (SGWB) sourced by the AD scalar field provides a new perspective probe by utilizing the naturally high-energy environment in the early cosmos when the AD mechanism operates. As the AD scalar condensate rolls toward its true vacuum and oscillates around it, it can lead to an era of rapid, non-perturbative particle production (e.g., Kitajima et al. (2018); Figueroa and Torrenti (2017); Lozanov and Amin (2019); Hiramatsu et al. (2021); Easther et al. (2007)), known as parametric resonance, and subsequent fragmentation, which can generate SGWB due to the large, time-dependent field inhomogeneities Friedberg et al. (1976); Coleman (1985); Kusenko (1997); Kusenko and Shaposhnikov (1998) produced therein. In the conventional SUSY-based framework, it was shown that parametric resonance of the AD condensates enabling baryogenesis cannot generate detectable GWs Kusenko and Mazumdar (2008); Kusenko et al. (2009); Zhou (2015) 111 In SUSY-based AD models, producing detectable strength of GWs can typically lead to $\eta_{B}\gg 10^{-10}$  Kusenko and Mazumdar (2008); Kusenko et al. (2009). Therefore,  Kusenko and Mazumdar (2008); Kusenko et al. (2009) considered GWs from AD flat-direction fields with $B-L=0$, which is not relevant for baryogenesis. Furthermore, recent lattice simulations revealed that the typical frequency of the GWs generated from SUSY AD models is higher than the currently detectable range Zhou (2015). Finally, the term $|\Phi|^{2}\log(|\Phi|/M_{\mathrm{Pl}})$ in the potential considered in Kusenko and Mazumdar (2008); Kusenko et al. (2009); Zhou (2015) exhibits a pathology at $\Phi=0$ (the impact of which on their GW results warrants re-examination), a pathology absent in our polynomial potential.. On the other hand, Q-balls, generically predicted in SUSY AD models, may result in a period of early matter domination (EMD) that induces GWs at second order through enhancing pre-existing primordial curvature perturbations. White et al. (2021); Kohri and Terada (2018). The Q-ball enhanced GW amplitude strongly depends on the speed of the transition from MD to RD White et al. (2021), with an upper limit given by the commonly used abrupt transition.
@@
The purpose of this work is two-fold: first, we present non-SUSY realizations of the AD mechanism that involve new physics below the electroweak scale and can operate in cosmic epochs long after inflation; second, demonstrate that parametric resonance sourced by AD condensates in such models naturally produces SGWB within reach of upcoming GW experiments. In particular, based on benchmark models with simple renormalizable scalar potentials, we show that in the viable parameter ranges for baryogenesis, parametric resonance of the AD condensate generates GWs with a peak frequency in the range of ${\cal O}(10-100)$ Hz, which can be explored by a combination of upcoming detectors, including Cosmic Explorer (CE) Abbott et al. (2017) and Einstein Telescope (ET) Hild et al. (2008), as well as DECIGO Kawamura et al. (2021) and Big Bang Observatory (BBO) Harry et al. (2006) (to capture the lower frequency tail).
@@
The first discovery of GWs by the LIGO-Virgo-KAGRA collaboration  Abbott et al. (2016a, b); Acernese et al. (2015) ushers in a new era of GW astronomy, followed by the more recent compelling evidence of SGWB identified by pulsar timing array (PTA) observations Carilli and Rawlings (2004); Weltman et al. (2020); Lentati et al. (2015); Babak et al. (2016); Agazie et al. (2023); Afzal et al. (2023); Antoniadis et al. (2023a, b). This study uncovers a generic class of well-motivated sources of SGWB, making a timely complement to the literature. The study also reveals a new potential synergy between GW detection and laboratory experiments at energy, intensity, and neutrino frontiers, such as LHC, DUNE, SHiP, FASER Aad et al. (2021); Aaboud et al. (2018); Bolton et al. (2020); Ariga et al. (2019); Ballett et al. (2020); Abe et al. (2019); Drewes et al. (2018); Barouki et al. (2022): interactions between the AD scalar $\Phi$ and the SM particles are necessary to transfer the asymmetry from the $\Phi$ field to SM baryons, while the characteristic new physics scale in the AD scalar sector is of ${\cal O}(0.1)$${\cal O}(10)$ GeV for viable models with detectable GW frequencies.
The rest of the Letter is organized as follows: in the next section, we present the benchmark models and the prediction for the baryon asymmetry; following that, we show the results of lattice simulation for GW signals from these models. Finally, we conclude with an outlook.
Benchmark Models.
I. AD scalars: background dynamics and asymmetry generation
@@
We consider a complex scalar field $\Phi$, which is a SM gauge singlet with canonical kinetic terms and Einstein gravity. We work with a simple renormalizable polynomial potential:
$V(\Phi)=\lambda_{\Phi}\left|\Phi\right|^{4}+m_{\Phi}^{2}\left|\Phi\right|^{2}-A(\Phi^{n}+\Phi_{\ast}^{n}),$ (1)
@@
with $n$ in principle being any integer $\leq 4$. We choose $n=2$ for concreteness; however, the results are not expected to change significantly with other choices for $n$. The first two terms manifest a $U(1)$ symmetry, which is broken by the third term as necessary to generate the $\Phi$-$\Phi^{\dagger}$ asymmetry, which later transfers to the baryon asymmetry. We consider the general possibility that $\Phi$, as a light spectator field during inflation, starts with a very large (even Planckian) field displacement Starobinsky and Yokoyama (1994); Cui et al. (2024); Hardwick et al. (2017); Markkanen et al. (2018, 2019), e.g., simply by initial condition, or by quantum fluctuation in the inflationary de Sitter space. We have taken into account the isocurvature constraints from CMB data, which is relevant to light spectator fields in general, and found that they can be generally satisfied (see Section .2 of the Supplemental Material for details). We denote the typical value at the end of inflation as $\Phi_{\mathrm{in}}$, the initial value for our calculations  [^2] Other mechanisms can be invoked to drive the $\Phi$ field to large values during inflation, like a non-minimal coupling to gravity or a direct coupling to the inflaton. Our results for the GWs and baryon asymmetry do not depend on the origin of $\Phi_{\mathrm{in}}$.. As long as the (radial) field is light, $H\gg H_{\mathrm{osc}}\simeq\sqrt{V_{,|\Phi|}/|\Phi|}$, where $H_{\mathrm{osc}}$ is the effective mass of $\Phi$ Cui et al. (2024), it remains frozen  [^3] Since we consider small $U(1)$ symmetry-breaking terms, the angular field will always be light when the radial field is light.. The initial phase angle of $\Phi$, $\theta=Arg(\Phi)$, is also set as $\theta_{\mathrm{in}}$ at the end of inflation in our observable patch of the universe. As the universe expands, the Hubble rate drops, and at $H\simeq H_{\mathrm{osc}}$ the field starts rolling. For our parameters, the rolling starts during the radiation-dominated (RD) era, well after the end of inflation.
@@
We analyze the motion of $\Phi$ in terms of its real and imaginary parts (as defined in $\Phi={1\over\sqrt{2}}(\phi_{R}+i\phi_{I})$), which obey the following equations of motion:
@@
where $m^{2}_{R,I}=m_{\Phi}^{2}\mp 2A$. The Hubble expansion rate in a RD Universe is $H\propto t^{1/2}$. The overdots denote derivatives with respect to cosmic time $t$. We see that the two components have different masses in the presence of $A\neq 0$, and thus, the trajectory in the $\phi_{R}$-$\phi_{I}$ plane is expected to resemble a shrinking ellipse instead of a straight line, even for vanishing initial velocities.
The background field evolution, neglecting the symmetry-breaking term ($A\ll m_{\Phi}^{2}$), is initially dominated by the $\left|\Phi\right|^{4}$ term when the amplitude scales as $\phi\propto 1/a$. When the amplitude drops below $|\Phi_{\ast}|=m_{\Phi}/\sqrt{\lambda_{\Phi}}$ at $t=t_{*}$, the $\left|\Phi\right|^{2}$ term dominates the oscillation and the fields evolve approximately as $\phi_{R,I}\sim a^{-3/2}\cos(m_{R,I}t+c_{R,I})$ (see Section .3 of the Supplemental Material for details). The $\Phi$-$\Phi^{\dagger}$ asymmetry, which will be transferred later to the baryon (or lepton) number, is $n_{\Phi}(t)=i\left(\Phi^{\dagger}\dot{\Phi}-\dot{\Phi}^{\dagger}\Phi\right)=\dot{\phi}_{\mathrm{R}}\phi_{{}_{\mathrm{I}}}-\phi_{{}_{\mathrm{R}}}\dot{\phi}_{\mathrm{I}}$. $n_{\Phi}$ directly corresponds to $n_{B}$ or $n_{L}$ in the absence of significant washout (see discussion later). Using the above approximations for $\phi_{R,I}$ and including the terms $\Gamma_{\Phi}\dot{\phi}_{R,I}$ to model the decay of $\Phi$ to SM states, we can write the co-moving baryon asymmetry generated during the time $t>t_{*}$ as Rubakov and Gorbunov (2017); Lloyd-Stubbs and McDonald (2021, 2023); Mohapatra and Okada (2021):
@@
[…]
where the exponential factor accounts for the decrease in $\Phi$ amplitude due to its decay to SM particles with a constant decay rate $\Gamma_{\Phi}$, enabling the asymmetry transfer to the SM baryons. The specific form of $\Gamma_{\Phi}$ depends on the decay channels, which we will elaborate on later. Evaluating the above integral for $t\rightarrow\infty$ and factoring in the expansion during the RD Universe, we express $n_{B}/s$ as
@@
$\frac{n_{B}}{s}=\begin{cases}\left(\frac{4\alpha^{3}}{\lambda_{\Phi}k_{T_{d}}^{6}}\right)^{1/4}\epsilon_{\Phi}\frac{m_{\Phi}M_{\mathrm{Pl}}}{T_{d}^{2}}\sin(2\theta);&\gamma_{\Phi}\gg 2\epsilon_{\Phi}\\ \left(\frac{\alpha^{3}k_{T_{d}}^{2}}{64\lambda_{\Phi}}\right)^{1/4}\frac{1}{\epsilon_{\Phi}}\frac{T_{d}^{2}}{m_{\Phi}M_{\mathrm{P}}}\sin(2\theta);&\gamma_{\Phi}\ll 2\epsilon_{\Phi}\end{cases}$ (5)
@@
where $M_{\mathrm{Pl}}\simeq 2.4\times 10^{18}$ GeV is the reduced Planck mass, and we define the dimensionless quantities: $\epsilon_{\Phi}=A/m_{\Phi}^{2}$, $\gamma_{\phi}=\Gamma_{\Phi}/m_{\Phi}$ and $k_{T_{d}}=\sqrt{\pi^{2}g(T_{d})/90}$. $T_{d}$ is the temperature of the Universe when the $\Phi$ decay is completed at $H(T_{d})=\Gamma_{\Phi}$, $\alpha$ is the initial fraction of energy density in the spectator sector $\Phi$ relative to the total energy density, and $\theta=\theta_{\mathrm{in}}$ is defined earlier. In case the asymmetry is first transferred to SM leptons, a redistribution factor of $28/79$ due to the EW sphaleron would be added in the RHS of Eq. (5).
@@
Eq. (5) is derived assuming that the asymmetry is generated when the potential is dominated by the $|\Phi|^{2}$ term; thus, we require that the background dynamics prior to $t_{*}$, defined as when $|\Phi(t_{\ast})|=m_{\Phi}/\sqrt{\lambda_{\Phi}}$, is not affected by the $A$ term. The corresponding condition on $A$ can be estimated as follows: The only term in the potential that involves the angular degree of freedom is the symmetry-breaking term $2A|\Phi(t)|^{2}\cos 2\theta(t)\subset V(|\Phi|,\theta)$, indicating an upper bound on the effective mass along the angular direction: $|m^{2}_{\theta}|\leq 2A$. The angular degree of freedom is frozen at $\theta_{\mathrm{in}}$ for $t<t_{*}$, if $|m^{2}_{\theta}|<H^{2}(t_{*})$ Lloyd-Stubbs and McDonald (2023), which translates to $\epsilon_{\Phi}\equiv A/m_{\Phi}^{2}<({\mu^{2}}/{16\alpha})\left({\phi_{\mathrm{in}}}^{2}/{M_{\mathrm{Pl}}}^{2}\right)$, where $\mu^{2}\equiv m_{\Phi}^{2}/(\lambda_{\Phi}\phi_{\mathrm{in}}^{2})$, and $\alpha$ is defined below Eq. 5. As we see in the next section, typically ${\phi_{\mathrm{in}}}\simeq M_{\mathrm{Pl}}$ and $\mu^{2}\lesssim 10^{-5}$ are required for detectable GW signals. With larger $A$, motion along the angular direction is non-negligible, and $\theta(t)$ would have already evolved towards its minimum before $t=t_{*}$, thus reducing $\sin\theta$ and suppressing the asymmetry (see Eq. (5)). For our benchmark examples, the chosen $A$ values satisfy this self-consistency condition.

### v2

# Probing baryogenesis with gravitational waves

@@ Abstract
Affleck-Dine (AD) baryogenesis is compelling yet challenging to probe because of the high-energy physics involved. We demonstrate that this mechanism can be generically realized with low-energy new physics without supersymmetry while producing detectable gravitational waves (GWs) sourced by parametric resonance of a light scalar field. In viable models, the scalar has a mass of $\mathcal{O}(0.1-10)$ GeV, yielding GWs with peak frequencies of $\mathcal{O}(10-100)$ Hz. This study further reveals a new complementarity between upcoming LIGO-frequency GW detectors and laboratory searches across frontiers of particle physics.
@@
Introduction. The observed excess of matter over antimatter in the universe, in terms of the baryon-to-photon ratio $\eta=\frac{n_{b}-n_{\bar{b}}}{n_{\gamma}}\approx 6\times 10^{-10}$ Aghanim et al. (2020), remains a long-standing puzzle in particle physics and cosmology. Affleck-Dine (AD) baryogenesis Affleck and Dine (1985) is a compelling solution to the puzzle, where the asymmetry emerges from the oscillation and decay of a baryon number-carrying complex scalar field. Originally formulated with supersymmetry (SUSY) Affleck and Dine (1985); Dine et al. (1996); Dine and Kusenko (2003); Enqvist and Mazumdar (2003), it involves flat direction moduli fields composed of superpartners of standard model (SM) particles. These light moduli fields can acquire large displacements during inflation, break C and $CP$ symmetries and subsequently roll down and oscillate to produce baryon asymmetry, while the inflaton still dominates the Universe Allahverdi and Mazumdar (2012) (alternatively, the AD field itself can be the inflaton Hertzberg and Karouby (2014); Cline et al. (2020a)).
Despite its appeal, experimentally testing the AD mechanism is challenging because of the high-scale physics involved in the inflationary era. Fortunately, the stochastic gravitational wave background (SGWB) generated by the oscillating AD field offers a promising new probe by utilizing the naturally high-energy environment in the early cosmos: oscillating scalar fields can trigger parametric resonance, i.e., rapid nonperturbative particle production, leading to large inhomogeneities and fragmentation that source GWs (e.g., Kitajima et al. (2018); Figueroa and Torrenti (2017); Lozanov and Amin (2019); Hiramatsu et al. (2021); Easther et al. (2007); Friedberg et al. (1976); Coleman (1985); Kusenko (1997); Kusenko and Shaposhnikov (1998)).
In conventional SUSY AD baryogenesis, parametric resonance does not generate detectable GWs Kusenko and Mazumdar (2008); Kusenko et al. (2009); Zhou (2015)   [^1] In SUSY AD models, producing detectable GWs can typically lead to $\eta_{B}\gg 10^{-10}$ Kusenko and Mazumdar (2008); Kusenko et al. (2009). Therefore, Kusenko and Mazumdar (2008); Kusenko et al. (2009) considered GWs from AD flat-direction fields with BL=0B\!-\!L=0, not relevant for baryogenesis. Furthermore, the potential term $|\Phi|^{2}\log(|\Phi|/M_{\mathrm{Pl}})$ considered in Kusenko and Mazumdar (2008); Kusenko et al. (2009); Zhou (2015) exhibits a pathology at $\Phi=0$, which is absent in our polynomial potential; its impact on GW emission warrants re-examination”. Meanwhile, SUSY AD models can lead to Q-balls, which may result in early matter domination and induce GWs at second order by enhancing preexisting curvature perturbations Kusenko and Mazumdar (2008); Kusenko et al. (2009). Q-ball-induced GWs strongly depend on the speed of the transition from matter to radiation domination White et al. (2021).
@@
The purpose of this work is twofold: first, we present non-SUSY realizations of the AD mechanism that involve new physics below the electroweak scale and can operate in cosmic epochs long after inflation; second, we demonstrate that parametric resonance sourced by AD condensates in such models naturally produces SGWB within reach of upcoming GW experiments.
@@
This study uncovers a generic class of well-motivated sources of SGWB and adds a timely complement to the rapidly growing field of probing particle physics with GWs, building on the ground-breaking discoveries of GWs by LIGO-Virgo-KAGRA collaboration Abbott et al. (2016a, b); Acernese et al. (2015) and the recent strong evidence for SGWBs revealed by pulsar timing arrays Carilli and Rawlings (2004); Weltman et al. (2020); Lentati et al. (2015); Babak et al. (2016); Agazie et al. (2023); Afzal et al. (2023); Antoniadis et al. (2023a, b). Based on simple benchmark models, we show that successful baryogenesis requires new physics in the AD sector with a characteristic scale $m_{\Phi}={\cal O}(0.1-10)$ GeV, and the resulting GWs have a peak frequency in the range of ${\cal{O}}(1-100)$ Hz, within reach of upcoming detectors, including the Cosmic Explorer (CE) Abbott et al. (2017) and Einstein Telescope (ET) Hild et al. (2008), as well as DECIGO Kawamura et al. (2021) and Big Bang Observatory (BBO) Harry et al. (2006) (for the low-frequency tail). The study further uncovers a new potential complementarity between GW detection and laboratory experiments at energy, intensity, and neutrino frontiers, such as the LHC, DUNE, SHiP, and FASER Aad et al. (2021); Aaboud et al. (2018); Bolton et al. (2020); Ariga et al. (2019); Ballett et al. (2020); Abe et al. (2019); Drewes et al. (2018); Barouki et al. (2022), in search for new physics, depending on the details of the asymmetry transfer interactions between the AD field and the SM.
Benchmark Models. I. AD scalars: background dynamics and asymmetry generation
@@
We consider a complex scalar field $\Phi$ (SM gauge singlet) with canonical kinetic terms, Einstein gravity, and a simple renormalizable polynomial potential:
$V(\Phi)=\lambda_{\Phi}\left|\Phi\right|^{4}+m_{\Phi}^{2}\left|\Phi\right|^{2}-A(\Phi^{n}+\Phi^{\ast n}),$ (1)
@@
where $n\leq 4$ is an integer. We choose $n=2$ for concreteness. Although other choices may lead to similar results, $n=2$ stands out as a simple and natural option   [^2] For example, with the alternative choice of $n=4$, the generated asymmetry quickly freezes in at a large value, requiring a very small value of A to reproduce the observable baryon asymmetryRubakov and Gorbunov (2017); Hertzberg and Karouby (2014). In contrast, with a quadratic symmetry-breaking term (n = 2), the oscillations of the $\Phi$-condensate around the origin continue to generate and cancel asymmetry over time, partially suppressing the initially large asymmetry . The first two terms manifest a $U(1)$ symmetry, which is broken by the third term as necessary to generate the $\Phi$-$\Phi^{\dagger}$ asymmetry, precursor to the baryon asymmetry. We consider the general possibility that $\Phi$, as a light spectator during inflation, starts with a large (even Planckian) field displacement Starobinsky and Yokoyama (1994); Cui et al. (2024); Hardwick et al. (2017); Markkanen et al. (2018, 2019), for instance, due to de Sitter fluctuations during inflation. We have verified that the isocurvature constraints from CMB data, which are relevant for any light spectator field, can be satisfied.We denote the typical value at the end of inflation as $\Phi_{\mathrm{in}}$, which is the initial value for our calculations   [^3] Other mechanisms can drive the $\Phi$ field to large values during inflation, like a non-minimal coupling to gravity or a direct coupling to the inflaton. Our results for GWs and $n_{b}-n_{\bar{b}}$ do not depend on the origin of $\Phi_{\mathrm{in}}$. . As long as the (radial) field is light, $H\gg H_{\mathrm{osc}}\simeq\sqrt{V_{,|\Phi|}/|\Phi|}$, where $H_{\mathrm{osc}}$ is the effective mass of $|\Phi|$ Cui et al. (2024), it remains frozen  [^4] Since we consider small $U(1)$ symmetry-breaking terms, the angular field will always be light when the radial field is light.. The initial phase of $\Phi$, $\theta=Arg(\Phi)$, is also set as $\theta_{\mathrm{in}}$ at the end of inflation in our observable patch of the Universe. As the Universe expands, the Hubble rate drops, and at $H\simeq H_{\mathrm{osc}}$ the field starts rolling. For our parameters, the rolling starts during the radiation-dominated (RD) era, well after the end of inflation.
@@
The real and imaginary parts of $\Phi\equiv\frac{1}{\sqrt{2}}(\phi_{R}+i\phi_{I})$ obey the following equations of motion:
@@
where $m^{2}_{R,I}=m_{\Phi}^{2}\mp 2A$. The Hubble expansion rate in an RD Universe is $H\propto t^{1/2}$. An overdot denotes derivative with respect to cosmic time $t$. If $A\neq 0$, the real and imaginary components have different masses, so their trajectories in the $\phi_{R}$-$\phi_{I}$ plane resemble shrinking ellipses (not straight lines), even for zero initial velocities.
The background field evolution (assuming $A\ll m_{\Phi}^{2}$) is initially dominated by the $\left|\Phi\right|^{4}$ term, where the amplitude scales as $\phi\propto 1/a$. When the amplitude drops below $|\Phi_{\ast}|=m_{\Phi}/\sqrt{\lambda_{\Phi}}$ at $t=t_{*}$, the $\left|\Phi\right|^{2}$ term dominates the oscillation and the fields evolve approximately as $\phi_{R,I}\sim a^{-3/2}\cos(m_{R,I}t+c_{R,I})$, with $c_{R,I}$ being a constant phase (see Supplemental Material, Section C, for details). The $\Phi$-$\Phi^{\dagger}$ asymmetry, which will be subsequently transferred to the baryon (or lepton) number, is $n_{\Phi}(t)\equiv i\left(\Phi^{\dagger}\dot{\Phi}-\dot{\Phi}^{\dagger}\Phi\right)=\dot{\phi}_{\mathrm{R}}\phi_{{}_{\mathrm{I}}}-\phi_{{}_{\mathrm{R}}}\dot{\phi}_{\mathrm{I}}$. In the absence of significant washout, $n_{\Phi}$ directly corresponds to $n_{B}$ or $n_{L}$ (see the Discussion later). Using the above approximations for $\phi_{R,I}$ and including the terms $\Gamma_{\Phi}\dot{\phi}_{R,I}$ to model the decay of $\Phi$ to SM states (the specific form of $\Gamma_{\Phi}$ depends on the decay channels, which we elaborate on later), the comoving baryon asymmetry generated during the time $t>t_{*}$ is Rubakov and Gorbunov (2017); Lloyd-Stubbs and McDonald (2021, 2023); Mohapatra and Okada (2021):
@@
$\int_{t_{*}}^{t}dt^{\prime}\cos{(m_{{}_{\mathrm{R}}}(t^{\prime}-t_{\ast}))}\sin{(m_{{}_{\mathrm{I}}}(t^{\prime}-t_{\ast}))}e^{-\Gamma_{\Phi}(t^{\prime}-t_{*})},$ (4)
which can be further simplified for $A\ll m_{\phi}^{2}$ (the detailed derivation is presented as Supplemental Material) and scales as
$\frac{a^{3}(t)}{a^{3}_{\mathrm{in}}}n_{B}(t)\propto\int_{t_{*}}^{t}dt^{\prime}\sin\left(\frac{2A}{m_{\Phi}}(t-t_{*})\right)e^{-\Gamma_{\Phi}(t-t_{*})}\,.$ (5)
Evaluating the above integral for $t\rightarrow\infty$ and factoring in the expansion during the RD Universe, we express $n_{B}/s$ as
@@
$\frac{n_{B}}{s}=\begin{cases}\left(\frac{4\alpha^{3}}{\lambda_{\Phi}k_{T_{d}}^{6}}\right)^{1/4}\epsilon_{\Phi}\frac{m_{\Phi}M_{\mathrm{Pl}}}{T_{d}^{2}}\sin(2\theta);&\gamma_{\Phi}\gg 2\epsilon_{\Phi}\\ \left(\frac{\alpha^{3}k_{T_{d}}^{2}}{64\lambda_{\Phi}}\right)^{1/4}\frac{1}{\epsilon_{\Phi}}\frac{T_{d}^{2}}{m_{\Phi}M_{\mathrm{Pl}}}\sin(2\theta);&\gamma_{\Phi}\ll 2\epsilon_{\Phi}\end{cases}$ (6)
@@
where $M_{\mathrm{Pl}}\simeq 2.4\times 10^{18}$ GeV is the reduced Planck mass, and we define the dimensionless quantities: $\epsilon_{\Phi}=A/m_{\Phi}^{2}$, $\gamma_{\phi}=\Gamma_{\Phi}/m_{\Phi}$ and $k_{T_{d}}=\sqrt{\pi^{2}g(T_{d})/90}$. $T_{d}$ is the temperature of the Universe characterizing $\Phi$ decay, as defined by $H(T_{d})=\Gamma_{\Phi}$, $\alpha$ is the initial fraction of energy density in the spectator sector $\Phi$ relative to the total energy density, and $\theta=\theta_{\mathrm{in}}$ is defined earlier. The exact form of $n_{B}/s$, from which Eq. (6) is derived, is given as Supplemental Material.
The change in the scaling of the final asymmetry with the normalized symmetry–breaking parameter $\epsilon_{\Phi}\equiv A/m_{\Phi}^{2}$ shown by the two branches of Eq. (6) can be understood simply in terms of the competition between two timescales: the oscillation period of the asymmetry and the lifetime of the scalar. The symmetry–breaking parameter $\epsilon_{\Phi}$ induces a mass splitting between the particle and antiparticle components of $\Phi$, leading to an oscillation of the charge asymmetry with frequency $\omega_{\Phi}\sim\Delta m_{\Phi}\sim\epsilon_{\Phi}m_{\Phi}$ and period $T_{\mathrm{asy}}\sim 1/(\epsilon_{\Phi}m_{\Phi})$. At the same time, the scalar decays with rate $\Gamma_{\Phi}$, corresponding to a lifetime $\tau_{\Phi}\sim 1/\Gamma_{\Phi}$. The net asymmetry transferred to the thermal bath is proportional to the time integral of an oscillatory source with an exponential envelope, schematically $n_{B}\propto\int_{0}^{\infty}dte^{-\Gamma_{\Phi}t}\sin(\omega_{\Phi}t)$, so the final result depends on whether the decay occurs before or after many oscillations. In the regime $\Gamma_{\Phi}\gg\omega_{\Phi}$ (i.e. $\tau_{\Phi}\ll T_{\mathrm{asy}}$), the integral is dominated by very early times, when the sine function has not yet completed a full oscillation and may be approximated as $\sin(\omega_{\Phi}t)\simeq\omega_{\Phi}t$. The generated asymmetry is then essentially fixed before sign–flipping oscillations set in, and scales linearly with the oscillation frequency (equivalently with the asymmetry parameter $\epsilon_{\Phi}$) as
$n_{B}\propto\int_{0}^{\sim 1/\Gamma_{\Phi}}dte^{-\Gamma_{\Phi}t}\omega_{\Phi}t\sim\frac{\omega_{\Phi}}{\Gamma_{\Phi}}\propto\epsilon_{\Phi}.$ (7)
By contrast, in the regime $\Gamma_{\Phi}\ll\omega_{\Phi}$ (i.e. $\tau_{\Phi}\gg T_{\mathrm{asy}}$), the source oscillates many times before the field decays, spending comparable time with positive and negative signs. As a result, contributions from successive oscillations partially cancel, and the net asymmetry is a residue arising from the slow decay of the envelope. To see the overall scale of the asymmetry for this limit, we note that the contribution from a single half-period of the oscillation is of order $\int dt\sin(\epsilon_{\Phi}m_{\Phi}t)\sim 1/(\epsilon_{\Phi}m_{\Phi})$, which sets the characteristic scale of the integral in Eq. (5). Including all subsequent oscillations leads to an alternating geometric series, which can be summed and consequently multiplies the result of the first oscillation by $1/2$ for the case with $\gamma_{\phi}/2\epsilon_{\Phi}\ll 1$. The final asymmetry thus scales as
$n_{B}\propto$ $\int_{0}^{\infty}dt\sin(\epsilon_{\Phi}m_{\Phi}t)$
$\simeq$ $\frac{1}{2}\times\int_{0}^{\pi/(2\epsilon_{\Phi}m_{\Phi}))}dt\sin(\epsilon_{\Phi}m_{\Phi}t)\sim\frac{1}{2}\frac{1}{\epsilon_{\Phi}m_{\Phi}}.$ (8)
In summary, a quadratic symmetry-breaking term induces a constant mass splitting between the real and imaginary components of $\Phi$, leading to a single well–defined oscillation frequency $\omega_{\Phi}\propto\epsilon_{\Phi}m_{\Phi}$. This simple harmonic structure makes the final asymmetry an explicit function of $\omega_{\Phi}$, and hence of $\epsilon_{\Phi}$, producing the characteristic two-regime behavior [^5] Note that the $\epsilon_{\Phi}\rightarrow 0$ limit is only reached through the upper branch of Eq. (6), leading to $n_{B}/s\rightarrow 0$ as expected. $n_{B}\propto\epsilon_{\Phi}$ for $\Gamma_{\Phi}\gg\omega_{\Phi}$ and $n_{B}\propto 1/\epsilon_{\Phi}$ for $\Gamma_{\Phi}\ll\omega_{\Phi}$, shown in Eq. (6). In the case where the asymmetry in the $\Phi$ field is first transferred to an SM lepton asymmetry $n_{L}$, which is later converted to a baryon asymmetry $n_{B}$ via the sphaleron process, the above result gets multiplied by a factor of $28/79$ since Harvey and Turner (1990); Asaka et al. (2002); Garcia-Bellido et al. (1999) $n_{B}/s=-(28/79)\times n_{L}/s\,.$
@@
Equation (6) is derived assuming that the asymmetry is generated when the potential is dominated by the $|\Phi|^{2}$ term, thus requiring that the background dynamics prior to $t_{*}$ is not affected by the $A$ term. The corresponding condition on $A$ is derived as follows: the only term in the potential that involves the angular degree of freedom is the symmetry-breaking term $2A|\Phi(t)|^{2}\cos 2\theta(t)\subset V(|\Phi|,\theta)$, indicating an upper bound on the effective mass along the angular direction: $|m^{2}_{\theta}|\leq 2A$. The angular degree of freedom is frozen at $\theta_{\mathrm{in}}$ for $t<t_{*}$ if $|m^{2}_{\theta}|<H^{2}(t_{*})$ Lloyd-Stubbs and McDonald (2023), which translates into $\epsilon_{\Phi}\equiv A/m_{\Phi}^{2}<({\mu^{2}}/{16\alpha})\left(\phi_{\mathrm{in}}^{2}/M_{\mathrm{Pl}}^{2}\right)$, where $\mu^{2}\equiv m_{\Phi}^{2}/(\lambda_{\Phi}\phi_{\mathrm{in}}^{2})$, and $\alpha$ is defined below Eq. 6. As we see in the next section, ${\phi_{\mathrm{in}}}\simeq M_{\mathrm{Pl}}$ and $\mu^{2}\lesssim 10^{-5}$ are required for detectable GW signals. With larger $A$, motion along the angular direction is non-negligible, and $\theta(t)$ would have already evolved toward its minimum before $t=t_{*}$, thus reducing $\sin\theta$ and suppressing the asymmetry [see Eq. (6)]. For our benchmark examples, the chosen $A$ values satisfy this self-consistency condition.